# A Transformer Model for Segmentation, Classification, and Caller Identification of Marmoset Vocalization

### v3

@@
A Transformer Model for Segmentation, Classification, and Caller Identification of Marmoset Vocalization
Bin Wu,1    Shinnosuke Takamichi,1,3    Sakriani Sakti,1,2    Satoshi Nakamura1,2,4
(1RIKEN Center for Advanced Intelligence Project (AIP), Japan
2Nara Institute of Science and Technology, Ikoma 630-0192, Japan
3Keio University, Keio 108-0073, Japan
4The Chinese University of Hong Kong, Shenzhen, China
1 Abstract
Marmoset, a highly vocalized primate, has become a popular animal model for studying social-communicative behavior and its underlying mechanism comparing with human infant linguistic developments. In the study of vocal communication, it is vital to know the caller identities, call contents, and vocal exchanges. Previous work of a CNN has achieved a joint model for call segmentation, classification, and caller identification for marmoset vocalizations. However, the CNN has limitations in modeling long-range acoustic patterns; the Transformer architecture that has been shown to outperform CNNs, utilizes the self-attention mechanism that efficiently segregates information parallelly over long distances and captures the global structure of marmoset vocalization. We propose using the Transformer to jointly segment and classify the marmoset calls and identify the callers for each vocalization.
2 Introduction
@@
Figure 1: The two-stream Transformer model. Two Transformer models process two spectral segments extracted from a dual-channel audio recording to map to a vocal call label. The call label of ’phee2’ denotes that the phee call comes from the second marmoset of the pair in the cage; the ’noise’ denotes the non-call segment between calls, used for segmenting calls to non-calls. We used two output heads to enable simultaneous call prediction, allowing detection of concurrent calls of the animal pair such as ’trill1’ and ’twitter2’.
3 Model
@@
4 Evaluation
We use F-score and accuracy to evaluate classification, segmentation, and caller identification with the same evaluation methods used in [9].
For caller evaluation, we convert the predicted labels into two separate segment files for a marmoset pair (e.g., ’tr2’ labels are converted to ’tr’ and added to the segment file of the second marmoset).We evaluate the two predicted segment files against corresponding annotated ones for the marmoset pair housed together.
@@ 4 Evaluation
5 Dataset and Experimental setup
5.1 Dataset
@@
We experimented on the dataset [7] designed to study close-range communication between marmosets. We used dual audio recordings from the dataset when two marmosets are together in a cage (excluding recordings when two marmosets are in separate cages). Each marmoset pair, familiar with each other after at least 3 months cohabitation before experiments, is placed in a homecage (77.5cm x 77.5cm x 147cm) with visual and auditory access. The home cage is located in an animal room with other cages positioned more than 90 cm away. The marmosets in other cages have no relation as a parent, siblings, or previous cage mates to the recorded marmosets.
The researchers collected vocalizations using light recorders (Panictech and Polend digital voice recorders, 6.9g), mounted with duct tape on a leather jacket, positioned close to the mouth ($<$5cm) of each marmoset of the recording pair (Upper subfigure (a) of Figure 3). The dataset includes 10 marmoset pairs housed together. The first five pairs (4 male-female pairs and 1 male-male pair) were recorded for about 1 to 2 hours a pair, with 8 to 12 sessions where each session lasted for 5 or 10 minutes. The second five pairs (1 male-female pair and 4 male-male pairs) were recorded for about 2 to 3 hours a pair, with 4 to 6 sessions where each session lasted for 30 minutes.
@@
The annotations of the dataset include the segment files for 10 marmoset pairs where each segment file of a marmoset contains a sequence of call labels with their beginning and end times. The 10 pairs include 8 pairs annotated by hand and 2 pairs annotated by a neural network and corrected by one observer. The manual annotations of the 8 pairs were split between two observers. Some sessions annotated by both observers have IOR (Inter-observer Reliability using Cohen’s Kappa) of 0.86 for call occurrence and IOR of 0.91 for the call type.
Table 1: Duration of training, development, and test sets.
@@
5.2 Data division
The dataset [7] consists of the dual audio recordings of 10 pairs of marmosets, aged from 1.5 to 10 years old, where the 5 female-male pairs are unrelated and the 5 male-male pairs are siblings. The annotations include 11 call types: trill, twitter, chirp, phee, trillphee, other, ek, tsik, chatter, peep, and infant cry (Figure 4). We divided the dataset with the data of pair1 as the test set, pair2 as the development set, and pair 3 to pair 10 as the training set with duration statistics (Table 1) and frequency statistics (Figure 4). All data come from annotated dual audio recordings of marmoset pairs housed together.
5.3 Experimental setup for systems
@@
5.3.1 Our CNN backbone model
We implemented our backbone CNN model with the same architecture as [9]. It consists of two CNN streams whose outputs are concatenated and linear-projected for the final prediction. Each CNN stream comprises 4 CNN modules. Each module comprises two identical convolutional layers followed by max-pooling. Whenever max-pooling is applied, the number of channels in the convolutional layers doubles in the subsequent module.
5.3.2 Proposed Transformer backbone model
We implemented our backbone Transformer model using Vision transformer that divides the original high-resolution $257\times 256$ linear spectral segments into $16\times 16$ patches. These patches are linearly embedded and then passed into a Transformer module with a model dimension of 384. The Transformer module consists of 6 blocks, each with 6-head self-attention and linear modules. The linear modules have a hidden dimension of 1536. The outputs of the two Transformer streams are concatenated and passed through a shared linear layer with a dimension of 1024 for the final prediction.
5.3.3 Proposed CNN/Transformer system with backbone model
@@
Table 2: Results of our CNN and proposed Transformer systems and [9]’s best system on the same dataset [7]
@@
Proposed Transformer 0.794 0.7576 0.8341
@@
6 Result and discussion
@@
Our CNN system outperforms the best system from [9] in both F-score and accuracy. Furthermore, the Transformer system surpasses the CNN system across all F-score and accuracy-related metrics.
The Transformer’s superior performance can be attributed to its ability to better capture overall patterns within each segment. However, the Vision Transformer is constrained by the patch-based resolution due to its quadratic complexity in input token length, resulting in lower model resolution compared to the CNN model. In future work, we aim to improve our Transformer system for better resolution modeling.
@@
7 Conclusion
We implemented a two-stream Transformer system that enables classification, segmentation, and caller identification for marmoset vocalization. This Transformer system outperforms previous CNN approaches. Our system can efficiently process long-hour or full-day marmoset recordings, providing rich information about vocal communication between marmosets in spontaneous interactions. This capability enhances the use of marmosets as an animal model for studying language evolution, development, and dysfunction in a communication context such as the comparative research between vocal development in marmosets and human infants in family environments.
8 Acknowledgements

### v4

# Learning Marmoset Vocal Patterns with a Masked Autoencoder for Robust Call Segmentation, Classification, and Caller Identification

@@
Abstract
The marmoset, a highly vocal primate, is a key model for studying social-communicative behavior. Unlike human speech, marmoset vocalizations are less structured, highly variable, and recorded in noisy, low-resource conditions. Learning marmoset communication requires joint call segmentation, classification, and caller identification—challenging domain tasks. Previous CNNs handle local patterns but struggle with long-range temporal structure. We applied Transformers using self-attention for global dependencies. However, Transformers show overfitting and instability on small, noisy annotated datasets. To address this, we pretrain Transformers with MAE—a self-supervised method reconstructing masked segments from hundreds of hours of unannotated marmoset recordings. The pretraining improved stability and generalization. Results show MAE-pretrained Transformers outperform CNNs, demonstrating modern self-supervised architectures effectively model low-resource non-human vocal communication.
Index Terms:
Animal call detection, marmoset vocalization, Transformer, self-supervised learning, MAE, ViT, segmentation, classification, caller identification.
I Introduction
@@
However, Transformer shows the problems of overfitting and unstable training on the target dataset. The available annotated two-stream dataset [15] specifically designed for caller annotation has a limited amount. We found that typical 12-layer Transformer overfits on the dataset just as CNN does. We have to decrease both the number of layers and model dimension to reduce overfitting. The complex task that involves segmentation, classification, and caller identification on noisy data also challenges Transformer training. We observed sudden drops in accuracy, some of which never recover to the original level. Such training instability is more severe when we use larger Transformer models.
To address the problem of the overfitting and unstable training of large Transformer with limited annotated data on challenging tasks, we propose using Masked Autoencoder (MAE) [19] to pretrain the Transformer. We applied MAE on hundreds of hours of marmoset recordings without any annotations. We found that after pretraining, the Transformer shows more stable training and almost no overfitting on the limited dataset. We frozed lower layers to extract MAE features and used higher layers for fine-tuning, enabling us to use a typical 12-layer Transformer without reducing model dimension. The MAE pretrained Transformer works best in practice.
@@
Figure 1: The two-stream Transformer model. The Transformer model can be initialized with MAE encoder pretrained by spectral segments without labels. We can freeze the first N layers for MAE feature extraction and the later N layers for fine-tuning. Two Transformer models process two spectral segments extracted from a dual-channel audio recording to map to a vocal call label. The call label of ’phee2’ denotes that the phee call comes from the second marmoset of the pair in the cage; the ’noise’ denotes the non-call segment between calls, used for segmenting calls to non-calls. We used two output heads to enable simultaneous call prediction, allowing detection of concurrent calls of the animal pair such as ’trill1’ and ’twitter2’.
II Model
@@
We can initialize each stream of Vision Transformer with identical parameters from the SSL model of MAE pretrained on spectral segments without labels. We froze the first M layers for MAE feature extraction and the later N layers for task-specific fine-tuning. This method retains the generalizability of MAE features while fine-tuning fewer parameters, making training faster and reducing overfitting risk for small annotated datasets.
@@
III Evaluation
We use F-score and accuracy to evaluate classification, segmentation, and caller identification, same as [16].
For caller evaluation, we convert predicted labels into two separate segment files for a marmoset pair (e.g., ’tr2’ labels are converted to ’tr’ and added to the segment file of the second marmoset).We evaluate two predicted segment files against corresponding annotated ones for marmoset pair housed together.
@@ III Evaluation
We also evaluate segmentation with the boundary F-score. The boundary F-score is the harmonic mean of boundary precision and recall, where precision is the number of correct intervals over the number of predicted intervals and recall is the number of correct intervals over the number of annotated intervals. A predicted interval is counted as correct only if there exists one annotated interval whose predicted begin and end times match the annotated begin and end times within a given threshold (which we used a tight threshold of 100 milliseconds compared to our model resolution of 50 milliseconds). Each predicted interval is only allowed to match one reference interval.
IV Dataset and Experimental setup
IV-A Dataset
@@
We experimented on the dataset [15] designed to study close-range communication between marmosets. We used dual audio recordings from the dataset when two marmosets are together in a cage (excluding recordings when two marmosets are in separate cages). The dataset includes 10 marmoset pairs housed together.
@@
TABLE I: Duration of training, development, and test sets of dual-stream annotated marmoset recordings [15] for evaluation of classification, segmentation, and caller identification.
@@
TABLE II: Duration of training, development, and test sets for MAE pretraining on marmoset recordings without annotations.
Data division Training set Development set Test set
Number of days 48 days 1 day 1 day
Recording time 639:38:19 17:45:57 18:38:12
IV-B Data division
The dataset [15] consists of the dual audio recordings of 10 pairs of marmosets, aged from 1.5 to 10 years old, where the 5 female-male pairs are unrelated and the 5 male-male pairs are siblings. The annotations include 11 call types: trill, twitter, chirp, phee, trillphee, other, ek, tsik, chatter, peep, and infant cry (Figure 4). We divided the dataset with the data of pair1 as the test set, pair2 as the development set, and pair 3 to pair 10 as the training set with duration statistics (Table I) and frequency statistics (Figure 4). All data come from annotated dual audio recordings of marmoset pairs housed together.
We collected longitudinal recordings for 50 days of a marmoset family with parents and one child in a sound-proof box. The recording is single-stream and we recorded more than 10 hours per day. We pretrained the MAE on the whole dataset (Table II) using 48 days for pretraining after roughly removing some empty audios and long silence periods without vocalization. The audios have no labels and we used MAE self-supervised learning on the marmoset linear spectrogram segments without annotations.
IV-C Experimental setup for systems
@@
IV-C1 Our CNN backbone model
We implemented our backbone CNN model with the same architecture as [16]. It consists of two CNN streams whose outputs are concatenated and linear-projected for the final prediction. Each CNN stream comprises 4 CNN modules. Each module comprises two identical convolutional layers followed by max-pooling. Whenever max-pooling is applied, number of channels in convolutional layers doubles in subsequent module.
IV-C2 Proposed Transformer backbone model
We implemented our backbone Transformer model using Vision transformer that divides the original high-resolution $257\times 256$ linear spectral segments into $16\times 16$ patches. The spectrogram patches reduce Transformer input lengths compared to pixels and balance the time and frequency resolutions. These patches are linearly embedded and then passed into a Transformer module with a model dimension of 384. The Transformer module consists of 6 blocks, each with 6-head self-attention and linear modules. The linear modules have a hidden dimension of 1536. The outputs of the two Transformer streams are concatenated and passed through a shared linear layer with a dimension of 1024 for the final prediction. We found that larger Transformer models (e.g., model dimension of 768 or deeper layers) would perform poorly on small amounts of available annotated data.
Figure 5: The accuracy of our CNN, proposed Transformer, and MAE Transformer on the development set during training epochs. The proposed MAE Transformer is the Transformer pretrained using a MAE model on marmoset recordings without labels. We observed that the MAE pretrained Transformer showed better training stability and less overfitting.
IV-C3 Proposed Transformer backbone model pretrained by MAE
We implemented our backbone Transformer model using an MAE encoder pretrained on the ImageNet training set [19] and 48 days of marmoset spectrogram segments (more than 10 hours per day). We used 12 layers of the pretrained Transformer, with the lower 6 layers frozen used for MAE feature extraction and the upper 6 layers for task-specific fine-tuning (where 6 layers performed the most stably and best among 3, 6, and 12 layers). We initialized identical parameters for the Transformers of both streams. We found that using a slightly larger model with model dimension 768 did not degrade performance on limited annotated data due to good pretrained initialization that stabilized training and MAE features that have good generalization power.
The pretrained MAE for marmoset recordings uses 12 layers, 12 attention heads, 768 embeddings with a masking ratio of 75%. The pretraining on the marmoset dataset included 400 training epochs with batch size 256, learning rate 0.00015, 40 warmup epochs, and 0.05 weight decay.
TABLE III: Separate classification, caller identification, and segmentation evaluations where segmentation evaluation uses the boundary F-score with 100ms tolerance. We used  [16]’s best system as a baseline and evaluated our CNN and proposed Transformer and MAE Transformer systems on the same dataset [15]. MAE_Transformer is the Transformer pretrained using a MAE model on marmoset recordings without labels.
Systems Call Acc. Caller Acc. Boundary F-score
Baseline 0.7212 0.7437 0.0180
CNN 0.7575 0.7866 0.0182
Transformer 0.7576 0.7867 0.0152
MAE_Transformer 0.7811 0.8119 0.0166
IV-C4 Proposed CNN/Transformer system with backbone model
@@
TABLE IV: Results of our CNN and proposed Transformer and MAE Transformer systems and [16]’s best system as a baseline on the same dataset [15]. The proposed MAE Transformer is the Transformer pretrained using a MAE model on marmoset recordings without labels.
@@
Proposed Transformer 0.7940 0.7576 0.8341
Proposed MAE Transformer 0.7998 0.7811 0.8195
@@
Proposed MAE Transformer 0.9909 0.9959 0.7811
V Result and discussion
@@
The Figure 5 shows real-time accuracy over 100 epochs on the development set during training of the CNN, Transformer, and MAE pretrained Transformer where MAE is pretrained with marmoset recordings without annotation. We observed that: 1) For CNN, the small model quickly converges on the limited training data within 25 epochs and then overfits without recovery. 2) For the Transformer model, overfitting occurs much later around 60 epochs, but its starting accuracy is low and shows a steep drop in accuracy due to unstable training. 3) For the MAE pretrained Transformer (6 frozen layers for MAE features and 6 layers for fine-tuning for MAE Transformer vs. 6-layer Transformer), the starting accuracy after the first epoch is even better than CNN. We did not observe any overfitting during fine-tuning, with accuracy continuing to increase until the end of 100 epochs. The MAE pretrained Transformer outperformed both the Transformer and CNN by a large margin.
Our proposed Transformer systems surpass the CNN system and baseline system [16] across all F-score and total accuracy (Table IV); the bootstrap one-tailed unpaired t-test on F-score comparing our CNN and MAE ViT is significant at the 0.05 level. Our proposed Transformer systems also achieve the best performance in separate evaluations of classification, and caller identification (Table III).
The Transformer performs best using global contextual modeling to better capture overall call patterns. However, the Vision Transformer patchizes input spectrograms, resulting in lower model resolution than the CNN model. The transformer prediction intervals are usually wider than annotated truth, penalized by strict segmentation evaluation such as 100ms tolerance boundary F-score (Table III); we also observed penalty at 50ms and 200ms tolerance boundary F-scores. In future work, we will improve Transformer system for better resolution modeling.
@@
VI Conclusion
We proposed self-supervised pretraining that enables robust marmoset call segmentation, classification, and caller identification with two-stream Transformer systems. The Transformer systems outperform previous CNN approaches. Our Transformer pretrained by MAE on hundreds of hours of marmoset recordings without annotations shows the best performance with more stable training and less overfitting compared to CNN and Transformer without pretraining. Our system efficiently processes long-hour or full-day marmoset recordings for vocal communication between marmosets in spontaneous interactions, advancing research on language evolution, development, and dysfunction.
VII Acknowledgements